# Spread them Apart: Towards Robust Watermarking of Generated Content

### v1

@@ 1 Introduction
Recent advances in generative models have brought the performance of image synthesis tasks to a whole new level. For example, the quality of the images generated by diffusion models (DMs, Croitoru et al. [2023], Rombach et al. [2022], Esser et al. [2024]) is now sometimes comparable to the one of the human-generated pictures or photographs. Compared to generative adversarial networks (GANs, Goodfellow et al. [2014], Brock et al. [2019], diffusion models allow the generation of high-resolution, naturally looking pictures and incorporate much more stable training, leading to a more diverse generation. More than that, the image generation process with diffusion models is more stable, controllable, and explainable. They are easy to use and are widely deployed as tools for data generation, image editing (Kawar et al. [2023], Yang et al. [2023]), music generation (Schneider et al. [2024]), text-to-image synthesis (Saharia et al. [2022], Zhang et al. [2023], Ruiz et al. [2023]) and in other multimodal settings.
Unfortunately, several ethical and legal issues may arise from the usage of diffusion models. On the one hand, since diffusion models can be used to generate fake content, for example, deepfakes (Zhao et al. [2021], Narayan et al. [2023]), it is crucial to develop automatic tools to verify that a particular digital asset is artificially generated. On the other hand, a dishonest user of the model protected by a copyright license can query it, receive the result of generation, and later claim exclusive copyright. In this work, we focus on the detection of the content generated by a particular model and the identification of the end-user who queried the model to generate a particular content. We develop a technique to embed the digital watermark into the generated content during the inference of the generative model, so it does not require retraining or fine-tuning the generative model. The approach allows not only to verify that the content was generated by a source model but also to identify the user who sent a corresponding query to the generative model. We prove that the watermark embedded is robust against additive perturbations of the content of a bounded magnitude.
@@
  • We apply the framework to watermark images generated by a diffusion model and prove that the watermark embedded is provably robust to the additive perturbations of a bounded magnitude that can be applied during the post-processing of the image.
  • Experimentally, we show that our approach outperforms competitors in terms of the robustness to different types of post-processing of the images aimed at watermark removal, such as brightness and contrast adjustment or gamma correction.
@@ 2 Related Work
2.1 Diffusion Model
Inspired by non-equilibrium statistical physics, Sohl-Dickstein et al. [2015] introduced the diffusion model to fit complex probability distributions. Ho et al. [2020] introduced a new class of models called Denoising Diffusion Probabilistic Models (DDPM) by establishing a novel connection between the diffusion model and the denoising scoring matching. Later, the Latent Diffusion Model (LDM) was developed to improve efficiency and reduce computational complexity, with the diffusion process happening within a latent space Rombach et al. [2022]. During training, the LDM uses an encoder $\mathcal{E}$ to map an input image $x$ to the latent space: $z=\mathcal{E}(x)$. For the reverse operation a decoder $\mathcal{D}$ is employed, so that $x=\mathcal{D}(z)$. During inference, the LDM starts with a noise vector $z\sim\mathcal{N}(0,I)$ in the latent space and iteratively denoises it. The decoder then maps the final latent representation back to the image space.
@@ 2.2 Watermarking of Digital Content
Watermarking has been recently adopted to protect the intellectual property of neural networks (Wu et al. [2020], Pautov et al. [2024]) and generated content (Kirchenbauer et al. [2023], Zhao et al. [2024], Fu et al. [2024]). In a nutshell, watermarking of generated content is done by injection of digital information within the generated image allowing the subsequent extraction. Existing methods of digital content watermarking can be divided into two categories: content-level watermarking and model-level watermarking. The methods of content-level watermarking operate in some representation of content, for example, in the frequency domain of the image signal (ó Ruanaidh et al. [1996], Cox et al. [1996]). When the image is manipulated in the frequency domain, the watermark embedding process can be adapted to produce watermarks that are robust to geometrical image transformations, such as rotations and translations (Wen et al. [2024]). Model-level watermarking approaches are designed to embed information during the generation process. In end-to-end methods, the models to embed and extract watermark are learned jointly (Zhu et al. [2018], Hayes and Danezis [2017]). In Yu et al. [2021], it was proposed to teach the watermark encoder on the training data of the generative model; such an approach yields a watermarking scheme that is conditioned on the generative model and its training dataset. This method was later adapted to latent diffusion models (Fernandez et al. [2023]) and unconditional diffusion models (Zhao et al. [2023]). In contrast, there are methods that do not require additional model training. These methods are designed to alter the output distribution of the generative model to embed previously learned watermark into the model or the content itself (Kirchenbauer et al. [2023], Wen et al. [2024]).
@@ 2.3 Robustness to Watermark Removal Attacks
Watermarking attacks are aimed at removing the watermark embedded into the model’s weights or generated content. In the prior works on removing the watermarks from generated images (Li et al. [2019], Cao et al. [2019]), the attack problem is formulated in terms of the image-to-image translation task, and methods to remove watermarks via an auxiliary generative adversarial network are presented. Other approaches (Hertz et al. [2019], Liang et al. [2021], Sun et al. [2023]) perform watermark removal in two steps: firstly, the visual watermark is localized within an image; secondly, it is removed via a multi-task learning framework.
In practice, watermarking scheme has to be robust to destructive and constructive attacks, or synthetic transformations of the data. Destructive transformations, such as brightness and contrast adjustment, geometric transformations, such as rotations and translations, compression methods, and additive noise are aimed at watermark removal by applying a transformation. In contrast, constructive attacks treat watermarks as noise and are aimed at the restoration of original content (Zhang et al. [2024]). It is usually done by applying purification techniques, such as Gaussian blur (Hosam [2019]) or image inpainting (Liu et al. [2021], Xu et al. [2017]).
Signal Processing Attacks focus on noise addition, compression, and filtering. Robust watermarking schemes based on frequency domain transformations and randomizing offered higher resilience against these types of attacks (Taran et al. [2019]).
@@ 3.1 Image Watermarking
In our approach, we focus on detection and attribution of the generated image simultaneously: while detection is aimed to verify whether a particular image is generated by a given model, attribution is aimed at determining the user who generated the image.
Suppose that we are given the generative model $f$ deployed in the black-box setting, i.e., as a service: in the generation phase a user $u_{i}\in[u_{1},\dots,u_{m}]$ sends a query to the model and receives a generated image $x\in\mathbb{R}^{d}$. If $x$ is a watermarked image, the owner of model $f$ should be able to identify that $x$ is generated by user $u_{i}$ by querying the model $f$. In our method, the image is watermarked during the generation phase, not during the post-processing. We formulate the process of watermarking and attribution in the following way:
@@
    When the watermarked object $x$ is received by the model owner, it extracts the watermark $w(u_{i}|x)$ using the secret $s(u_{i})$ of the user $u_{i}$ and compares it with the watermark $w(u_{i})$ assigned to the user $u_{i}$. Following the previous works (Yu et al. [2021], Fernandez et al. [2023]), we compute the bitwise distance $d(w(u_{i}|x),w(u_{i}))$ between $w(u_{i}|x)$ and $w(u_{i})$:
@@
    where $\tau_{1}\ll n$ and $\tau_{2}\gg 0.$ This procedure is known as the double-tail detection (Jiang et al. [2023]).
@@ 3.2 The Probability of Incorrect Attribution
We assume that the watermark $w(u_{i})$ attributed to the user $u_{i}$ is drawn randomly and uniformly from the set of all possible $n-$bit watermarks, $\{0,1\}^{n}$. Following the prior works (Fernandez et al. [2023]), we formulate the detection problem as the hypothesis test. In case of a single user $u_{i}$, we define the null hypothesis $\mathcal{H}_{0}=$ “the object $x$ is generated not by $u_{i}$” and the alternative hypothesis $\mathcal{H}_{1}=$ “the object $x$ is generated by $u_{i}$”. Additionally, under the null hypothesis, we assume that the $j^{\prime}$th bit in the watermark $w(u_{i}|x)$ extracted from $x$ is the same as the $j^{\prime}$th bit from $w(u_{i})$ with the probability $p_{i}$.
@@
$FPR(m)\leq\mathbb{P}_{w^{\prime}\sim\{0,1\}^{n}}[\exists u_{j}\in[u_{1},\dots,u_{m}]:$ (4)
$d(w^{\prime},w(u_{j}))\in[0,\tau_{1}]\cup[\tau_{2},n]]\leq$
$\sum_{u_{j}\in[u_{1},\dots,u_{m}]}FPR(1)|_{u_{j}}=\hat{p}.$
@@
Remark.
In our experiments, the probability $p_{i}$ from above is estimated to be close to $\frac{1}{2}.$
@@ 3.3 Robustness to Watermark Removal Attacks
When the user $u_{i}$ receives the watermarked image $x$, it can post-process it to obtain the other image, $x^{\prime}$, which does retain the sufficient part of the watermark $w(u_{i})$. The transition from $x$ to $x^{\prime}$ may be done by applying an image transformation, such as brightness or contrast adjustment, Gaussian blur, or additive noise. The other approach is to perform an adversarial attack on the generative model to erase the watermark (Jiang et al. [2024]). In our settings, we assume that the generative model is deployed as the black-box service with limited access to the API, so an adversary can not apply white-box adversarial attacks (Jiang et al. [2023]).
4 Method
In this section, we provide a detailed description of the proposed approach, its implementation details, and the robustness guarantee against additive watermarking removal attacks of bounded magnitude.
@@ 4.1 Embedding and Extraction of the Watermark
Suppose that $f$ is the generative model. Recall that the user $u_{i}\in[u_{1},\dots,u_{m}]$ receives a pair $(w(u_{i}),s(u_{i}))$ after the registration in the service, where both the watermark and the secret are unknown to the user and are privately kept by the owner of $f$. Let $x$ be the generated image. Then, the watermark embedding process is described as follows:
@@
    where $x_{j}$ is the intensity of the $j^{\prime}$th pixel of $x$. To increase the robustness to watermark removal attacks, we apply additional regularization to $x$:
    $\min_{j\in[1,\dots,n]}|x_{a_{j}}-x_{b_{j}}|\geq\epsilon,$ (7)
    where $\epsilon>0$ is the scalar parameter.
To perform detection and attribution of the given image $x$, the owner of the generative model firstly constructs $m$ watermarks $w(u_{1}|x),\dots,w(u_{m}|x)$ by reversing the implication from the Eq. 6. Namely, given the secret $s(u_{i})=\{a_{1},\dots,a_{n},b_{1},\dots,b_{n}\}$ of user $u_{i}$, the watermark bits are restored by the following rule:
$\begin{cases}x_{a_{j}}\geq x_{b_{j}}\implies w(u_{i}|x)_{j}=0,\\ x_{a_{j}}<x_{b_{j}}\implies w(u_{i}|x)_{j}=1.\par\end{cases}$ (8)
Remark.
@@
where $\xi_{i}$ is the indicator function from the Eq. 5. Note that if $\xi_{i}=0$ for all $\ i\in[1,\dots,m],$ then $x$ is identified as image not generated by $f$.
@@ 4.2 Implementation Details
In this subsection, we describe the watermarking procedure. First of all, we have to note that in the Stable Diffusion model, the latent vector $z$ produced by the U-Net is then decoded back into the image space using a VAE decoder: $x=\mathcal{D}(z)$. To embed the watermark into an image, we optimize a special two-component loss function with respect to the latent vector $z.$ The overall loss is written as follows:

### v2

# arXiv:2502.07845v2[cs.CV] 27 Feb 2026

@@ 1 Introduction
Recent advances in generative models have brought the performance of image synthesis tasks to a whole new level. For example, the quality of the images generated by diffusion models [croitoru2023diffusion, rombach2022high, esser2024scaling] is now comparable to that of human-drawn pictures or photographs. Compared to generative adversarial networks [goodfellow2014generative, brock2019large], diffusion models allow the generation of high-resolution, diverse, and naturally looking pictures. More than that, the generation process with diffusion models is more stable, controllable, and explainable. They are easy to use and are widely deployed as tools for data generation, image editing [kawar2023imagic, yang2023paint], music generation [schneider2024mousai], text-to-image synthesis [saharia2022photorealistic, zhang2023adding, ruiz2023dreambooth], and in other multimodal settings.
Unfortunately, several ethical and legal issues may arise from the usage of diffusion models. On the one hand, since diffusion models can be used to generate fake content, for example, deepfakes [zhao2021multi, narayan2023df], it is crucial to develop automatic tools to verify that a particular digital object is artificially generated. On the other hand, a dishonest user of the model protected by a copyright license can query it, receive the result of generation, and later claim exclusive copyright, possibly disobeying the copyright agreement. In this work, we focus on the detection of the images generated by a particular diffusion model and on the identification of the end-user who queried the model to generate an image. We develop a technique to embed the digital watermarks into the generated content during the inference of the generative model. Namely, during generation, we optimize the latent representation of the image so that the output image satisfies the set of predefined end-user-specific inequalities. These inequalities are later used to reliably identify the end-user, so the method solves both content detection and user attribution tasks. The detailed description of the approach is presented in the upcoming sections.
@@
  • We apply the framework to watermark images generated by a diffusion model and prove that the watermarks embedded are guaranteed to be robust to the additive perturbations of a bounded magnitude, multiplicative perturbations, and exponentiations.
  • Experimentally, we show that our approach yields watermarks that are robust to different types of post-processing of the images aimed at watermark removal, such as brightness and contrast adjustment, gamma correction, or a white-box adversarial attack.
@@ 2 Related Work
2.1 Diffusion Models
Inspired by non-equilibrium statistical physics, [sohldickstein2015deepunsupervisedlearningusing] introduced the diffusion model to fit complex probability distributions. Later, [ho2020denoisingdiffusionprobabilisticmodels] introduced a new class of models called Denoising Diffusion Probabilistic Models (DDPM) by establishing a novel connection between the diffusion model and the denoising score matching. Shortly after, the Latent Diffusion Model (LDM) was developed to improve efficiency and reduce computational complexity, with the diffusion process conducted in a latent space [rombach2022high]. During training, the LDM uses an encoder $\mathcal{E}$ to map an input image $x$ to the latent space, $z=\mathcal{E}(x)$. For the reverse operation, a decoder $\mathcal{D}$ is employed, so that $\hat{x}=\mathcal{D}(z)$. During inference, the LDM initializes a noise vector $z\sim\mathcal{N}(0,I)$ in the latent space and iteratively denoises it. The decoder then maps the final latent representation back to the image space.
@@ 2.2 Watermarking of Digital Content
Watermarking has been recently adopted to protect the intellectual property of neural networks [wu2020watermarking, pautov2024probabilistically] and generated content [kirchenbauer2023watermark, zhaoprovable, fu2024watermarking]. In a nutshell, watermarking of generated content is done by injecting digital information within the generated content, allowing the subsequent extraction. Existing methods of digital content watermarking can be divided into two categories: content-level watermarking and model-level watermarking. The methods of content-level watermarking operate within a particular representation of content, for example, in the frequency domain of the image signal [o1996watermarking, cox1996secure]. Notably, when the image is manipulated in the frequency domain, the watermark embedding process can be adapted to produce watermarks that are robust to geometrical image transformations, such as rotations and translations [wen2024tree]. In contrast, model-level watermarking approaches are designed to embed information during the generation process. In end-to-end methods, the models to embed and extract watermark are trained jointly [zhu2018hidden, hayes2017generating]. In [yu2021artificial], it was proposed to teach the watermark encoder on the training data of the generative model; such an approach yields a watermarking scheme that is conditioned on the generative model and its training dataset. This method was later adapted to latent diffusion models [fernandez2023stable] and unconditional diffusion models [zhao2023recipe]. In contrast, there are methods that do not require additional model training. These methods are designed to alter the output distribution of the generative model to embed previously learned watermark into the model or the content itself [kirchenbauer2023watermark, wen2024tree].
@@ 2.3 Robustness to Watermark Removal Attacks
Watermark removal attacks are aimed at removing the watermark embedded into the model’s weights or generated content. In the prior works on removing the watermarks from generated images [li2019towards, cao2019generative], the attack problem is formulated in terms of the image-to-image translation task, and methods to remove watermarks via an auxiliary generative adversarial network are presented. Other approaches [hertz2019blind, liang2021visible, sun2023denet] perform watermark removal in two steps: firstly, the visual watermark is localized within an image; secondly, it is removed via a multi-task learning framework. In practice, a watermarking scheme has to be robust to destructive and constructive attacks, or synthetic transformations of the data. Destructive transformations, such as brightness and contrast adjustment, geometric transformations, such as rotations and translations, compression methods, and additive noise, are aimed at watermark removal by applying a transformation. In contrast, constructive attacks treat watermarks as noise and are aimed at the restoration of original content [zhang2024robust]. It is usually done by applying purification techniques, such as Gaussian blur [hosam2019attacking] or image inpainting [liu2021wdnet, xu2017automatic].
@@ 3.1 Image Watermarking
In our approach, we focus on detection and attribution of the generated image simultaneously: while detection is aimed at verifying whether a particular image is generated by a given model, attribution is aimed at determining the user who generated the image.
Suppose that we are given the generative model $f$ deployed in the black-box setting, i.e., as a service: in the generation phase, a user $u_{i}\in[u_{1},\dots,u_{m}]$ sends a query to the model and receives a generated image $x\in\mathbb{R}^{d}$. If $x$ is a watermarked image, the owner of model $f$ should be able to identify that $x$ is generated by user $u_{i}$ by querying the model $f$. In our method, the image is watermarked during the generation phase, not during the post-processing. We formulate the process of watermarking and attribution in the following way:
@@
    When the watermarked object $x$ is received by the model owner, it extracts the watermark $w(u_{i}|x)$ using the secret $s(u_{i})$ of the user $u_{i}$ and compares it with the watermark $w(u_{i})$ assigned to the user $u_{i}$. To do so, we compute the bitwise distance $d(w(u_{i}|x),w(u_{i}))$ between $w(u_{i}|x)$ and $w(u_{i})$:
@@
where $\tau_{1}\ll n$ and $\tau_{2}\gg 0.$ This procedure is known as the double-tail detection [jiang2023evading] and is effective against watermark flipping attacks.
@@ 3.2 The Probability of Incorrect Attribution
We assume that the watermark $w(u_{i})$ attributed to the user $u_{i}$ is drawn randomly and uniformly from the set of all possible $n$-bit watermarks, $\{0,1\}^{n}$. We formulate the detection problem as a hypothesis test. In case of a single user $u_{i}$, we define the null hypothesis $\mathcal{H}_{0}=$ “the object $x$ is generated not by $u_{i}$” and the alternative hypothesis $\mathcal{H}_{1}=$ “the object $x$ is generated by $u_{i}$”. Additionally, under the null hypothesis, we assume that the $j^{\prime}$th bit in the watermark $w(u_{i}|x)$ extracted from $x$ is the same as the $j^{\prime}$th bit from $w(u_{i})$ with the probability $p_{i}$.
@@
$FPR(m)$ $\leq\ \mathbb{P}_{w^{\prime}\sim\{0,1\}^{n}}[\exists u_{j}\in[u_{1},\dots,u_{m}]:d(w^{\prime},w(u_{j}))\in[0,\tau_{1}]\cup[\tau_{2},n]]\leq$ (4)
$\leq\sum_{u_{j}\in[u_{1},\dots,u_{m}]}FPR(1)|_{u_{j}}=\hat{p}$.
@@
Remark 2
In our experiments, the probabilities $p_{i}$ from above are estimated to be close to $\frac{1}{2}.$
@@ 3.3 Robustness to Watermark Removal Attacks
When the user $u_{i}$ receives the watermarked image $x$, they can post-process it to obtain the other image, $x^{\prime}$, which does not retain a sufficient part of the watermark $w(u_{i})$. The transition from $x$ to $x^{\prime}$ may be done by applying an image transformation, such as brightness or contrast adjustment, Gaussian blur, or additive noise. The other approach is to perform an adversarial attack on the generative model to erase the watermark [jiang2024watermark].
4 Spread them Apart: Pixel level
In this section, we provide a detailed description of the proposed approach, its implementation details, and the robustness guarantee against additive removal attacks of bounded magnitude. Note that in this section, we describe the embedding of the watermarks on the pixel level; an extension of the method is discussed in the subsequent sections.
@@ 4.1 Embedding and Extraction of the Watermark
Suppose that $f$ is the generative model. Recall that the user $u_{i}\in[u_{1},\dots,u_{m}]$ is assigned a pair $(w(u_{i}),s(u_{i}))$ after the registration in the service, where both the watermark and the secret are unknown to the user and are privately kept by the owner of $f$. Let $x$ be the generated image. Then, the watermark injection and reconstruction are done as follows:
@@
    where $x_{j}$ is the intensity of the $j^{\prime}$th pixel of $x$. To increase the robustness to watermark removal attacks, given $\varepsilon>0$, we apply additional regularization to $x$:
    $\min_{j\in[1,\dots,n]}|x_{a_{j}}-x_{b_{j}}|\geq\epsilon.$ (7)
To perform detection and attribution of the given image $x$, the owner of the generative model firstly constructs $m$ watermarks $w(u_{1}|x),\dots,w(u_{m}|x)$ by reversing the implication from the Eq. 6. Namely, given the secret
$s(u_{i})=\{a_{1},\dots,a_{n},b_{1},\dots,b_{n}\}$
of user $u_{i}$, the watermark bits are restored by the following rule:
$\begin{cases}x_{a_{j}}\geq x_{b_{j}}\implies w(u_{i}|x)_{j}=0,\\ x_{a_{j}}<x_{b_{j}}\implies w(u_{i}|x)_{j}=1.\end{cases}$ (8)
Remark 3
@@
where $\xi_{i}$ is the indicator function from the Eq. 5. Note that if $\xi_{i}=0$ for all $i\in[1,\dots,m],$ then $x$ is identified as image not generated by $f$. If the argument in Eq. 9 is not unique, the algorithm abstains from attribution.
@@ 4.2 Implementation Details
In this subsection, we describe the watermarking procedure. In our settings, the generative model $f$ is Stable Diffusion [rombach2022high]. First of all, we have to note that in the Stable Diffusion model, the latent vector $z$ produced by the UNet is then decoded back into the image space using a VAE decoder, $\mathcal{D}$. To embed the watermark into an image, we optimize a special two-component loss function with respect to the latent vector $z.$ The overall loss is written as follows:
@@
$\mathcal{L}_{wm}=\sum_{i=1}^{n}\min((-1)^{w_{i}}(x_{a_{i}}-x_{b_{i}})+\varepsilon,0),\quad x=\mathcal{D}(z).$ (11)
Here, $\varepsilon$ is the watermark robustness threshold. Note that the larger the value of $\varepsilon$ is, the more robust the watermark is to additive perturbations. The second term, $\mathcal{L}_{qual}$, is introduced to preserve the quality of the generated image. The value $\mathcal{L}_{qual}$ is a difference in image quality measured by the LPIPS metric [zhang2018perceptual], which acts as a regularization.
The optimization is performed over $T=700$ steps of the Adam optimizer with the learning rate of $8\times 10^{-3}$, where every $100$ iteration, the learning rate is halved. When the convergence is reached, the ordinary Stable Diffusion post-processing of the image is performed. The coefficients $\lambda_{wm}$ and $\lambda_{qual}$ are determined experimentally and set to be $0.9$ and $150$, respectively, the value of $\varepsilon$ was set to be $\varepsilon=0.2$. Schematically, the process of watermark embedding and extraction is presented in Figure 1.
4.3 Robustness Guarantee
The watermark embedded by our method is robust against additive watermark removal attacks of a bounded magnitude. Namely, let the watermark $w(u_{i}|x)$ be embedded in $x$ with the use of the secret $s(u_{i})=\{a_{1},\dots,a_{n},b_{1},\dots,b_{n}\}$ of the user $u_{i}$. Let $\Delta_{i}=\frac{|x_{a_{i}}-x_{b_{i}}|}{2}$. Then, the following lemma holds.
Lemma 1
Let $\varepsilon\in\mathbb{R}^{d}$ and $\Delta_{i_{1}}\leq\Delta_{i_{2}}\leq\dots\leq\Delta_{i_{n}}$. Then, if $\|\varepsilon\|_{\infty}<\Delta_{i_{k}}$, then $d(w(u_{i}|x+\varepsilon),w(u_{i}|x))<k.$
The proof is presented in the supplementary material. This lemma provides a lower bound on the $l_{\infty}$ norm of the additive perturbation $\varepsilon$ applied to $x$, which is able to erase at least $k$ bits of the watermark $w(u_{i}|x))$ embedded into $x$.
5 Spread them Apart: Several Watermarks Instead of One
Here, we discuss an extension of the proposed method to provide robustness to other types of input perturbations, such as rotations and translations. An extension is based on a simultaneous embedding of the watermark in the pixel space and special functions, which are invariant to certain transformations. An intuition behind this extension is as follows. Suppose that the transform $\gamma:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is invariant under parametric perturbation $\phi:\mathbb{R}^{d}\times\Theta\to\mathbb{R}^{d}$, namely, $\gamma(x)=\gamma(\phi(x,\theta))$ for all $x\in\mathbb{R}^{d},\ \theta\in\Theta$. Then, if the watermark is successfully embedded into $\gamma(x)$, it becomes robust under perturbation $\phi$.
5.1 Invariants in the Frequency Domain
where $\Theta$ is the set of parameters of perturbation $\phi$. In this work, we use two invariants discussed in [lin1993towards], formulated as theorems below.
Theorem 5.1
Let $h(x,y)$ be an integrable nonnegative function, and its Fourier transform
$H(\omega_{x},\omega_{y})=\int_{\infty}^{\infty}\int_{-\infty}^{\infty}h(x,y)e^{-i(x\omega_{x}+y\omega_{y})}dxdy=A(\omega_{x},\omega_{y})e^{-i\psi(\omega_{x},\omega_{y})}$ (13)
be twice differentiable. Then the function $A(\omega_{x},\omega_{y})$ is translation invariant.
Theorem 5.2
$\tilde{H}(\omega,k)=\int_{-\infty}^{\infty}\int_{0}^{2\pi}\tilde{h}(r,t)e^{-i(kt+\omega r)}dtdr=\tilde{A}(\omega,k)e^{-i\tilde{\psi}(\omega,k)},$ (14)
where $\tilde{A}(\omega,k)$ is the magnitude and $\tilde{\psi}(\omega,k)$ is the phase. If $\tilde{h}(r,t)$ is an integrable nonnegative function and its Fourier-Mellin transform $\tilde{H}(\omega,k)$ is twice differentiable, then the function $\tilde{A}(\omega,k)$ is invariant under rotation.
Remark 4
We refer to invariants from Theorems 5.1-5.2 as to $\gamma_{t}$ and $\gamma_{r}$, respectively.
5.2 Three Watermarks Instead of One